Using the below information: Sales revenue $937,600 Less: cost of good sold Fixed 257,205 Variable 171,278 Gross profits $509,117 Less: operating expenses Fixed 202,533 Variable 54,369 Operating profits $252,215 Less: interest expense 29,066 Net profit before taxes $223,149 Less: taxes (rate 25%) 55,787 Net profits after taxes $167,362 Allen Products LP, wants to do a scenario analysis for the coming year. The pessimistic prediction for sales is $894,000; the most likely amount of sales is $1,124,000; and the optimistic prediction is $1,270,000. Allen's income statement for the most recent year is shown above, a. Use the percent-of-sales method, the income statement for December 31, 2019 and the sales revenue estimates to developpessimistic, most likely, and optimistic pro forma income statements for the coming year. b. Explain how this method could result in overstatement of profits for the pessimistic case and understatement of profits for the most likely and optimistic cases. c. Restate the pro foma income statement in prepared to part a. to incorporate assumptions about 2019 costs $257,205 of the cost of goods sold is fixed, the rest is variable. 202,533 of the operating expense is fixed rest is variable. All interest is fixed

Pro Forma Income Statement Allen Products, Inc. for the Year Ended December 31, 2020 Sales: Less: Cost of goods sold:_____ Gross profits: ___ / _____% Less: Operating expense:______/____% Operating profits:_______ Less: Interest expense:_____/______% Net profits before taxes:_______ Taxes (25%):_______ Net profits after taxes:________
Answer #1

Part A

Pro Forma Income Statement

Allen Products, Inc.

for the Year Ended December 31, 2020

Pessimistic

Sales

894000

Less: cost of goods sold

408558

45.7%

Gross profits

485442

Less: operating expenses

244956

27.4%

Operating profits

240486

Less: interest expense

27714

3.1%

Net profit before taxes

212772

Taxes (25%)

53193

Net profit after taxes

$159579

Cost of goods sold = (257205+171278)/937600 = 45.7%

Operating expenses = (202533+54369)/937600 = 27.4%

Interest expense = 29066/937600 = 3.1%

Pro Forma Income Statement

Allen Products, Inc.

for the Year Ended December 31, 2020

Most Likely

Sales

1124000

Less: cost of goods sold

513668

45.7%

Gross profits

610332

Less: operating expenses

307976

27.4%

Operating profits

302356

Less: interest expense

34844

3.1%

Net profit before taxes

267512

Taxes (25%)

66878

Net profit after taxes

$200634

Pro Forma Income Statement

Allen Products, Inc.

for the Year Ended December 31, 2020

Optimistic

Sales

1270000

Less: cost of goods sold

580390

45.7%

Gross profits

689610

Less: operating expenses

347980

27.4%

Operating profits

341630

Less: interest expense

39370

3.1%

Net profit before taxes

302260

Taxes (25%)

75565

Net profit after taxes

$226695

Part B

According to the simple percent-of-sales method it is assumed that all costs are variable. In reality environment there are some of the expenses which are fixed in nature. In the pessimistic case, there is decrease in all costs due to decrease in sales level but in real environment there is decrease in fixed costs. On the contrary, all costs increases in case of optimistic but in real situation only variable costs increases. In case of pessimistic case there is understatement of costs and overstatement of profits whereas the opposite happens in the case of optimistic scenario.

Part C

Pro Forma Income Statement

Allen Products, Inc.

for the Year Ended December 31, 2020

Pessimistic

Sales

894000

Less: cost of goods sold

Fixed

257205

Variable

163602

18.30%

Gross profits

473193

Less: operating expenses

Fixed

202533

Variable

51852

5.80%

Operating profits

218808

Less: interest expense

29066

Net profit before taxes

189742

Taxes (25%)

47436

Net profit after taxes

$142306

Cost of goods sold = 171278/937600 = 18.3%

Operating expenses = 54369/937600 = 5.8%

Pro Forma Income Statement

Allen Products, Inc.

for the Year Ended December 31, 2020

Most Likely

Sales

1124000

Less: cost of goods sold

Fixed

257205

Variable

205692

18.30%

Gross profits

661103

Less: operating expenses

Fixed

202533

Variable

65192

5.80%

Operating profits

393378

Less: interest expense

29066

Net profit before taxes

364312

Taxes (25%)

91078

Net profit after taxes

$273234

Pro Forma Income Statement

Allen Products, Inc.

for the Year Ended December 31, 2020

Optimistic

Sales

1270000

Less: cost of goods sold

Fixed

257205

Variable

232410

18.30%

Gross profits

780385

Less: operating expenses

Fixed

202533

Variable

73660

5.80%

Operating profits

504192

Less: interest expense

29066

Net profit before taxes

475126

Taxes (25%)

118782

Net profit after taxes

$356344